# Wave-U-Mamba: An End-To-End Framework For High-Quality And Efficient Speech Super Resolution

### v2

@@ C Results
Fig. 3 illustrates the sample mel spectrograms in training our model. We can see the checkerboard artifacts in Fig. 3c, but by iterative training, artifacts are mitigated and high-frequency patterns are captured more precisely. This is consistent with the previous finding [36] that long training could lead to the vanishing of checkerboard artifacts in many up-sampling applications.
@@
Table 1: Comparison of Log-Spectral Distance (LSD) on various input sampling rates with target sampling rate of 48 kHz with other baseline models.
@@
Objective Performance Comparison We have conducted an objective comparison of the model performance with other State-Of-The-Art (SOTA) models using LSD. Table C shows that our model Wave-U-Mamba excels other models in terms of LSD, meaning that our model effectively captures high-frequency patterns. Additionally, our model exhibits a relatively consistent performance for audio chunks of any sampling rate. We can infer that this is due to the robustness of waveform representations whereas for other models input mel representations can have varying amounts of active pixels.
Ablation studies As presented in table C, we conducted a series of experiments with various design configurations of the Wave-U-Mamba model, and then compared the LSD to better understand the functioning of some of the key components of the model. Wave-U-Mamba-deep refers to the same model architecture with the number of DownsampleBlock and UpsampleBlock increased to 5, which is set to 4 on default in Fig. 1. The dimension size at the bottleneck layer is also doubled. The resulting LSD of Wave-U-Mamba-deep is consistent with the idea that projecting LR signal into a relatively low-dimensional space is sufficient for accurate reconstruction. We also explored the impact of incorporating the L1 loss in the waveform domain in training. We observe that it introduced metallic artifacts in the reconstructed waveform. Lastly, removing MambaBlock from the model resulted in a marked degradation in performance, implying that Mamba is a crucial component of our model given that it functions as a global dependency extractor.
Subjective Performance Comparison As seen from the table C, MOS results suggest that our model has the best perceptual quality among all the other baseline models in improving LR speech signal to HR speech signal. Especially, our results have improved MOS scores of LR speech with a margin of 1.44.
Table 2: Comparison of Mean Opinion Score (MOS), inference time, and the number of parameters with other baseline models. The evaluation of ground truth High-Resolution (HR) signal and Low-Resolution (LR) signal have also been conducted to show the margin of improvement in the perceptual quality that Speech Super-Resolution (SSR) models achieve.
Model MOS $\uparrow$ Inference Speed (ms/s) $\downarrow$ # Parameters (M) $\downarrow$
@@
Table 3: Ablations of design choices of Wave-U-Mamba with regard to Log Spectral Distance (LSD) on various input sampling rates with target sampling rate of 48 kHz.
@@
Efficiency Comparison To compare the efficiencies of the models, we compared model size in gigabytes (GB), mean inference speed, and the number of parameters as criteria. To calculate the inference speed, we first randomly sample files from VCTK, then estimate the mean inference time in milliseconds for processing one second of input speech using Monte-Carlo simulation with 50 iterations. Inference speed is measured using a single A100 GPU. As indicated in Table C, our model contains fewer than 2% of the parameters of the baseline models and exhibits an inference speech 536.3 times faster than AudioSR.
@@ V Conclusion
In this work, we present Wave-U-Mamba, a fully end-to-end speech super-resolution model that operates in the waveform domain. The proposed model employs a U-structured architecture with generative adversarial networks for training. Wave-U-Mamba demonstrates the state-of-the-art performance on VCTK-test dataset, surpassing existing methods without reliance on additional models or pretrained weights. Furthermore, the model is capable of upsampling speech at any input resolution, from a sampling rate of 4 kHz to 24 kHz, to a high-resolution output at 48 kHz, consistently delivering high-quality results. The model’s efficiency in upsampling low-resolution speech signals is mainly attributed to its simple architecture, allowing for both computational efficiency and superior performance.

### v3

@@ IV-C Results
Fig. 3 illustrates the sample mel spectrograms in training our model. We can see the checkerboard artifacts in Fig. 3c, but by iterative training, artifacts are mitigated and high-frequency patterns are captured more precisely. This is consistent with the previous finding [38] that long training could lead to the vanishing of checkerboard artifacts in many upsampling applications.
@@
TABLE I: Comparison of Log-Spectral Distance (LSD) on various input sampling rates with target sampling rate of 48 kHz with other baseline models.
@@
Objective Performance Comparison We have conducted an objective comparison of model performance with other State-Of-The-Art (SOTA) models using LSD. Table IV-C shows that our model Wave-U-Mamba excels other models in terms of LSD, meaning that our model effectively captures high-frequency patterns. Additionally, our model exhibits a relatively consistent performance for audio chunks of any sampling rate. We can infer that this is due to the robustness of waveform representations, whereas for other models input mel representations can have varying amounts of active pixels.
Subjective Performance Comparison As seen in the table II, the result of the MOS suggest that our model has the best perceptual quality among all the other baseline models in improving LR speech signal to HR speech signal. In particular, our results have improved MOS scores of LR speech with a margin of 1.44.
Ablation studies As presented in table III, we conducted a series of experiments with various design configurations of the Wave-U-Mamba model and then compared the LSD to better understand the functioning of some of the key components of the model. Wave-U-Mamba-deep refers to the same model architecture with the number of DownsampleBlock and UpsampleBlock increased to 5, which is set to 4 by default in Fig. 1. The dimension size at the bottleneck layer is also doubled. The resulting LSD of Wave-U-Mamba-deep is consistent with the idea that projecting LR signal into a relatively low-dimensional space is sufficient for accurate reconstruction. We also explored the impact of incorporating L1 loss in the waveform domain in training. We observe that it not only deteriorated the performance but also introduced metallic artifacts in the reconstructed waveform. Lastly, removing MambaBlock from the model resulted in a marked degradation in performance, implying that Mamba is a crucial component of our model given that it functions as a global dependency extractor.
TABLE II: Comparison of Mean Opinion Score (MOS), inference time, and the number of parameters with other baseline models. The evaluation of ground truth High-Resolution (HR) and Low-Resolution (LR) signals is included to demonstrate the margin of perceptual quality improvement achieved by Speech Super-Resolution (SSR) models.
Model MOS $\uparrow$ Inference time (ms) $\downarrow$ # Parameters (M) $\downarrow$
@@
TABLE III: Ablations of design choices of Wave-U-Mamba with regard to Log Spectral Distance (LSD) on various input sampling rates with target sampling rate of 48 kHz.
@@
Efficiency Comparison To compare the efficiencies of the models, we used mean time to generate a second of speech fragment and the number of parameters as criteria. To calculate the inference speed, we first randomly sample files from VCTK, then estimate the mean inference time in milliseconds for generating one second of HR speech using a Monte Carlo simulation with 50 iterations. Inference speed is measured using a single A100 GPU. As indicated in Table II, our model contains fewer than 2% of the parameters of the baseline models and exhibits an inference speech 536.3 times faster than AudioSR.
@@ V Conclusion
In this work, we present Wave-U-Mamba, a fully end-to-end speech super-resolution model that operates in the waveform domain. The proposed model employs a U-structured architecture with generative adversarial networks for training. Wave-U-Mamba demonstrates the state-of-the-art performance on VCTK-test dataset, surpassing existing methods without relying on additional models or pretrained weights. Furthermore, the model is capable of upsampling speech at any input resolution, from a sampling rate of 4 to 24 kHz, to a high-resolution output at 48 kHz, consistently delivering high-quality results. The model’s efficiency in upsampling low-resolution speech signals is mainly attributed to its simple architecture, allowing for both computational efficiency and superior performance.
VI Acknowledgements
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT)(IITP-2025-RS-2024-00436857); Institute of Information & communications Technology Planning & Evaluation (IITP) grantfunded by the Korea government(MNIST)(No. RS-2019-||190079.Artificial Intelligence Graduate School Program(Korea University))